OpenStudy (anonymous):

Compare each pair of numbers use <. =. > 2/45. 1/30

OpenStudy (anonymous):

Not exactly orthodox but... 2/45 = 1/22.5. Which is bigger out of 22.5 and 30? And a bigger denominator with the same numerator means the number is...?

OpenStudy (anonymous):

the denom are 30 and 45 30x45=1350 (2/45)(30/30)= 60/1350 (1/30)(45/45)=45/1350 therefore 60/1350 > 45/1350 also 2/45 > 1/30 ans...